6. What is [H3O+] in a solution of 0.075M HNO2 and 0.030M NaNO, Ka of HNO2 = 4.5x10-5
Answer #1

We know that, Gaven, [salt] - [ NaNO₂] = 0·030M [Acid] = [HNO,] = 0·075M ka = 4:58 105 pka = - log ka = - log (4.5x105) = - [T-- --- - TEATT- TTTTTTT 1 1 1 1 1 . L. 2 4 -1- - - pH = 4-34+ (-1-52) + (-1-12) 1 - 1 - 0 1 1 - 24K | 1 - | | | | | | PAT-TE